# Gravitational waves from cosmic strings in LISA: reconstruction pipeline and physics interpretation

### v1

@@ Appendix B Templates for Non-Standard pre-BBN Cosmologies
In this appendix, we briefly outline how the templates for non-standard pre-BBN cosmologies, introduced in section 4.3, have been constructed. Let us start by considering the cases in which there are modifications to the equation of state of the early universe before the onset of radiation domination or the strings are created during an inflationary era. In both situations, the spectrum in the fundamental mode is of the form in eq. (10). For $f<f_{\rm rd}$, the spectra remains unmodified and thus the standard expression may be used. For $f>f_{\rm rd}$, however, the full SGWB, including all the harmonic modes of emission of loops, is given by
@@
$=$ $\frac{\Omega_{\rm rd}}{\zeta(q)}\left(\frac{f_{\rm rd}}{f}\right)^{d}\mathcal{H}^{(q-d)}_{N_{\rm rd}}+\Omega_{\rm gw}(f,q,G\mu,\alpha)-\Gamma\sum_{j=1}^{N_{\rm rd}}\frac{j^{-q}}{\zeta(q)}{\Omega_{\rm gw}}^{1}\,,$
where $N_{\rm rd}=f/f_{\rm rd}$ and the last term may be computed as described in Appendix A after substituting $N$ by $N_{\rm rd}$.
@@
Appendix C Models with non-gravitational radiation
The main text studies string models where the only channel for string decay is into GWs. Outside of these minimally-coupled strings, several models exist which have other, sometimes dominant, modes of energy loss.
Consider the work of refs. Hindmarsh:2021mnl ; Hindmarsh:2022awe , which suggests modeling an Abelian–Higgs (AH) network of strings by multiplying a Nambu–Goto (NG) SGWB by a parameter $f_{\text{NG}}$, representing the non-gravitational radiation:
Interestingly, the uncertainty in the case of $\log_{10}(G\mu)=-12.0$, $f_{\text{NG}}=0.1$ is less than the other points with $f_{\text{NG}}=0.1$, but its absolute error is greater when compared to the points adjacent to it. The reconstructions of this point prefer heavier strings with lower Nambu–Goto fraction than the injected values. While it is not immediately clear why this is the case, the answer may lie in which parts of the string SGWB are in the LISA window. Much of the structure in the string SGWB is in the radiation-to-matter-transition ‘bump’ and in the changing-degree-of-freedom ‘steps’. If the portion of the SGWB in the LISA window lacks much of this structure (as one might get by simply lowering a high-tension SGWB), then a slightly heavier but less-radiative string might be able to provide a better fit by SGWBinner’s metrics.

### v2

@@ Appendix B Templates for Non-Standard pre-BBN Cosmologies
In this appendix, we briefly outline how we have constructed the templates for non-standard pre-BBN cosmologies, introduced in section 4.3. Let us start by considering the cases in which there are modifications to the equation of state of the early universe before the onset of radiation domination described in section 2.4.1 (or the strings are created during an inflationary era in section 2.4.2). In this case, we resort to Model I to develop the templates. As shown in Cui:2017ufi , the spectrum in the fundamental mode is of the form in eq. (10). As in Appendix A, we then need to sum over the contributions of all relevant harmonic modes of emission to obtain a complete template for the signal. It is straightforward to show that, for $f<f_{\rm rd}$, the spectra is unaffected by the change in the cosmological background and thus the template for model I may be used. For $f>f_{\rm rd}$, however, the full SGWB, including all the harmonic modes of emission of loops, is given by
@@
$=$ $\frac{\Omega_{\rm rd}}{\zeta(q)}\left(\frac{f_{\rm rd}}{f}\right)^{d}\mathcal{H}^{(q-d)}_{N_{\rm rd}}+\Omega_{\rm gw}(f,q,G\mu,\alpha)-\Gamma\sum_{j=1}^{N_{\rm rd}}\frac{j^{-q}}{\zeta(q)}{\Omega_{\rm gw}}^{1}(f/j,G\mu,\alpha)\,,$
where $N_{\rm rd}=f/f_{\rm rd}$ and the last term may be computed as described in Appendix A after substituting $N$ by $N_{\rm rd}$. In the last equality, we introduced the definition of generalized harmonic number in the first term and used the fact that $\sum_{j=N_{\rm rd}}^{+\infty}=\sum_{j=1}^{+\infty}-\sum_{j=1}^{N_{\rm rd}}$ and (34) in the last term.